# SpaFL: Communication-Efficient Federated Learning with Sparse Models and Low Computational Overhead

### v1

@@ Abstract
The large communication and computation overhead of federated learning (FL) is one of the main challenges facing its practical deployment over resource-constrained clients and systems. In this work, SpaFL: a communication-efficient FL framework is proposed to optimize sparse model structures with low computational overhead. In SpaFL, a trainable threshold is defined for each filter/neuron to prune its all connected parameters, thereby leading to structured sparsity. To optimize the pruning process itself, only thresholds are communicated between a server and clients instead of parameters, thereby learning how to prune. Further, global thresholds are used to update model parameters by extracting aggregated parameter importance. The generalization bound of SpaFL is also derived, thereby proving key insights on the relation between sparsity and performance. Experimental results show that SpaFL improves accuracy while requiring much less communication and computing resources compared to sparse baselines.
@@ 1 Introduction
FL is a distributed machine learning framework in which clients collaborate to train a machine learning (ML) model without sharing private data [1]. In federated learning (FL), clients perform multiple epochs of local training using their own datasets and communicate model updates with a server. Different from a standard centralized setting, FL systems are typically deployed on edge devices such as mobile or Internet of Things (IoT) devices, which have limited computing and communication resources. However, current ML models are typically too large and complex to be trained and deployed for inference by edge devices. Moreover, large model sizes can induce significant FL communication costs on both devices and communication networks. Hence, the practical deployment of FL over resource-constrained devices and systems requires optimized computation and communication costs for both edge devices and communication networks. This has motivated lines of research focused on reducing communication overhead in FL [2; 3], training sparse models in FL [4; 5; 6; 7; 8; 9], and optimizing model architectures to find a compact model for inference [10; 11; 12]. The works in [2; 3] proposed training algorithms such as quantization, gradient compression, and transmitting the subset of models in order to reduce the communication costs of FL. However, the associated computational overhead of these existing algorithms remains high since devices have to train a dense model. In [4; 5; 6; 7; 8; 9], FL algorithms in which devices train and communicate sparse models are proposed. However, works in [4; 5] used unstructured pruning, which is difficult to gain the computation efficiency in practice. Moreover, the computation and communication overhead can still be large if model sparsity is not high. In [6; 7; 8; 9], although they investigated the structured sparsity, they fixed channel sparsity patterns for clients or did not optimize the pruning process. Furthermore, the FL approaches of [10; 11; 12] can significantly increase computation resource usage by training multiple models for resource-constrained devices. Clearly, despite a surge of literature on sparsity in FL, there is still a need to develop new FL algorithms that can find sparse model structures with optimized communication efficiency and low computational overhead to operate on resource-constrained devices.
@@ 2.2 Training and Finding Sparse Models in FL
To reduce the computation and communication overhead of complex ML models during training, the idea of embedding FL algorithms with pruning has emerged. In [4; 5; 6; 7; 8; 9; 17; 18; 19; 20; 21; 22; 23; 24; 25], the clients train sparse models and communicate sparse model parameters to reduce computation and communication overhead. To improve the aggregation phase with sparse models, the works in [17; 20; 21] perform averaging only between overlapping parameters to avoid information dilution by excluding zero value parameters. The work in [18] obtained a sparse model by selecting a particular client to prune an initial dense model and then performed training in a similar way to FedAvg. In [4; 24], the authors presented binary masks adjustment strategy to improve the performance of sparse models and communication efficiency. The work in [25] progressively pruned a dense model for sparsification and analyzed its convergence. In [19; 22], the clients optimized personalized sparse models by exchanging lottery tickets [26] at every communication round. The work in [5] obtained personalized sparse models by $l_{1}$ norms constraints and the correlation between local and global models. In [8], the authors proposed dual pruning scheme for both local and global models to reduce the communication costs. The FL framework of [23] allows clients to train personalized sparse models in a decentralized setting without a central server. Although these works [17; 18; 4; 24; 25; 19; 22; 5; 23] adopted sparse models during training, they used unstructured pruning, which is difficult to improve the computation efficiency in practice. Meanwhile, with structured sparsity, the authors [7] proposed a training scheme that allows clients to train smaller submodels of a global model. In [9], clients train set of submodels with fixed channel sparsity patterns depending on their computing capabilities. The work in [6] studied structured sparsity by adjusting clients’ channel activation probabilities. However, the works in [7; 9] fixed sparsity patterns and did not optimize sparse model structures. Although [6] optimized channel activation probabilities, the communication cost of downlink still remains high as a server broadcasts whole parameters. Similar to our work, in [27; 28], only binary masks are communicated and optimized by training auxiliary variables to learn sparse model structures. However, the work in [27] approximated binarization step using a sigmoid function during forward propagation. In [28], the downlink communication costs remained the same as that of FedAvg. In [10; 11; 29], clients perform neural-architecture-search by training multiple models to find optimized and sparse models to improve computational and memory efficiency at inference phase. However, in practice, clients often have limited resources to support the computationally intensive architecture search process [30]. Therefore, most prior works either adopted unstructured pruning or they still required extensive computing and communication costs for finding optimal sparse models. In contrast to the prior art, in the proposed SpaFL framework, we find sparse model structures with structured sparsity by optimizing and communicating trainable thresholds for filter/neurons.
@@ 5.2 Baselines
We compare SpaFL with multiple state of the art baselines that studied sparse model structures in FL. In FedAvg [1], every client trains a global dense model and communicates whole model parameters. FedPM [28] trains and communicates a binary mask while freezing model parameters. In HeteroFL [7], each client trains and communicates $p$-reduced models, which remove the last $1-p$ output channels in each layer. In Fjord [9], each client randomly samples a model from a set of $p$-reduced models, which drops out $p\%$ of filter/neurons in each layer. Local only performs local training with the introduced pruning method without any communications. For the sparse FL baselines, the average target sparsity is set to 0.5 following the configurations in [28; 7; 9].
@@ 5.3 Main Results
In Table LABEL:tab:1 and Fig. 2, we present the averaged accuracy, communication costs, number of FLOPs during training, and convergence rate for each algorithm. We consider all uplink and downlink communications to calculate the communication cost of each algorithm. We also provide the details of the FLOPs measure in the Supplementary document. We average the model densities of SpaFL when a model achieved the best accuracy during training. From these results, we observe that SpaFL outperforms all baselines while using the least amount of communication costs and number of FLOPs. The achieved model densities are 35.36%, 30.57%, and 35.38%, for FMNIST, CIFAR-10, and CIFAR-100, respectively. We also observe that SpaFL uses less resources and performs better than HetroFL and Fjord, which deployed structured sparse models across clients. Although FedPM reduced uplink communication costs by communicating only binary masks, its downlink cost is the same as FedAvg. In SpaFL, since the clients and the server only exchange thresholds, we can significantly reduce the communication costs compared to baselines that exchange the subset of model parameters such as HeteroFL and Fjord. Moreover, SpaFL significantly achieved better performance than Local, which did not communicate trained thresholds. Local achieved 51.2%, 50.1%, and 53.6% model densities for each dataset, respectively. We can see that communicating trained thresholds can make models sparser and achieve better performance. This also corroborates the analysis of Theorem 1. Hence, SpaFL can efficiently improve model performance with small computation and communication costs. In Fig. 2, we show the convergence rate of each algorithm. We can see that the accuracy of SpaFL decreases and then keeps increasing. The initial accuracy drop is from pruning while global thresholds are not trained enough. As thresholds keep being trained and communicated, clients learn how to prune their model, thereby gradually improving the performance with less active filter/neurons.
@@ 6 Conclusion
One major limitation is that SpaFL cannot explicitly control the sparsity of clients. Since we enforce sparsity through the regularizer term, we need to run multiple experiments to find values for desired sparsity. Another limitation is that our analysis requires a bounded loss function. Meanwhile, in practice, most loss functions may admit bounds that have a large value. For broader impact, SpaFL can reduce not only the computation and communication costs of FL training, but also those of inference phase due to sparsity. Hence, SpaFL can improve the sustainability of FL deployments, and more broadly, of AI.
@@ A.1 Implementation Detail
We run all experiments on NVIDIA A100 GPUs with PyTorch. In Table LABEL:tab:model, we provide detailed information of model architectures for each dataset. For the FMNIST dataset, we use the Lenet-5-Caffe model, which is Caffe variant of Lenet-5. The Lenet model has 430500 of model parameters and 580 of trainable thresholds. For the CIFAR-10 dataset, we use a CNN model of seven layers used in Zhou et al. [2019]. It has 807366 of model parameters and 1418 of trainable thresholds. The ResNet-18 model is adopted for the CIFAR-100 dataset with 11159232 of model parameters and 4800 of thresholds. We use a stochastic gradient optimizer with momentum of 0.9. For FMNIST with the Lenet model, we use $\eta(t)=0.001$, $E=5$, a batch size of 64, and $\alpha=0.002$. For CIFAR-10, we use $\eta(t)=0.01$, $E=5$, a batch size of 16, and $\alpha=0.00015$. For CIFAR-100, we use $\eta(t)=0.01$, $E=7$ decayed by $0.993$ at each communication round, a batch size of 64, and $\alpha=0.0007$. All trainable thresholds are initialized to zero. We noticed that too large sparsity coefficient $\alpha$ can dominate the training loss, resulting in masking whole parameters in a certain layer. Following the implementation of Liu et al. [2020], if a certain layer’s density becomes less than 1%, the corresponding trainable thresholds will be reset to zero to avoid masking whole parameters.
Table 4: Model architectures used in our experiments
@@ A.1.1 More details about baselines
We compare SpaFL with sparse baselines that investigated structured sparsity. In FedAvg McMahan et al. [2017], every client trains a global dense model and communicates whole model parameters. We used the equal weighted average for the model aggregation. FedPM Isik et al. [2022] optimizes a binary mask while freezing model parameters. Clients only transmit their arithmetically coded binary masks to the server, and the server broadcasts real-valued probability masks to the clients. We use Adam optimizer with learning rate of $0.1$ as done in Isik et al. [2022]. HeteroFL Diao et al. [2021] selects $\lceil pC\rceil$ channels of each layer, where $0\leq p\leq 1$ and $C$ is the number of channels, to make $p$ reduced submodels. Clients train and communicate $p$ reduced submodels during training. We set $p=0.5$ following Diao et al. [2021]. Fjord Horvath et al. [2021] samples $p$ from a uniform distribution $\mathcal{U}(p_{\text{min}},p_{\text{max}})$. After sampling $p$, clients train $p$ reduced submodel by selecting the first $\lceil pC\rceil$ channels of each layer. We set $p_{\text{min}}=0.4$ and $\boldsymbol{p}_{\text{max}}=0.6$ Horvath et al. [2021]. We provide the learning rates of the baselines in the following table.

### v2

@@ Abstract
The large communication and computation overhead of federated learning (FL) is one of the main challenges facing its practical deployment over resource-constrained clients and systems. In this work, SpaFL: a communication-efficient FL framework is proposed to optimize sparse model structures with low computational overhead. In SpaFL, a trainable threshold is defined for each filter/neuron to prune its all connected parameters, thereby leading to structured sparsity. To optimize the pruning process itself, only thresholds are communicated between a server and clients instead of parameters, thereby learning how to prune. Further, global thresholds are used to update model parameters by extracting aggregated parameter importance. The generalization bound of SpaFL is also derived, thereby proving key insights on the relation between sparsity and performance. Experimental results show that SpaFL improves accuracy while requiring much less communication and computing resources compared to sparse baselines. The code is available at https://github.com/news-vt/SpaFL_NeruIPS_2024
@@ 1 Introduction
FL is a distributed machine learning framework in which clients collaborate to train a machine learning (ML) model without sharing private data [1]. In federated learning (FL), clients perform multiple epochs of local training using their own datasets and communicate model updates with a server. Different from a classical, centralized ML, FL systems are typically deployed on edge devices such as mobile or Internet of Things (IoT) devices, which have limited computing and communication resources. However, current ML models are typically too large and complex to be trained and deployed for inference by edge devices. Moreover, large model sizes can induce significant FL communication costs on both devices and communication networks. Hence, the practical deployment of FL over resource-constrained devices and systems requires optimized computation and communication costs for both edge devices and communication networks. This has motivated lines of research focused on reducing communication overhead in FL [2; 3], training sparse models in FL [4; 5; 6; 7; 8; 9], and optimizing model architectures to find a compact model for inference [10; 11; 12]. The works in [2; 3] proposed training algorithms such as quantization, gradient compression, and transmitting the subset of models in order to reduce the communication costs of FL. However, the associated computational overhead of these existing algorithms remains high since devices have to train a dense model. In [4; 5; 6; 7; 8; 9], FL algorithms in which devices train and communicate sparse models are proposed. However, the works in [4; 5] used unstructured pruning, which is difficult to gain the computation efficiency in practice. Moreover, the computation and communication overhead can still be large if model sparsity is not high. In [6; 7; 8; 9], the authors investigated the structured sparsity, however, the solutions therein either fixed the channel sparsity patterns for clients or did not optimize the pruning process. Furthermore, the FL approaches of [10; 11; 12] can significantly increase computation resource usage by training multiple models for resource-constrained devices. Clearly, despite a surge of literature on sparsity in FL, there is still a need to develop new FL algorithms that can find sparse model structures with optimized communication efficiency and low computational overhead to operate on resource-constrained devices.
@@ 2.2 Training and Finding Sparse Models in FL
To reduce the computation and communication overhead of complex ML models during training, the idea of embedding FL algorithms with pruning has recently attracted attention. In [4; 5; 6; 7; 8; 9; 17; 18; 19; 20; 21; 22; 23; 24; 25], the clients train sparse models and communicate sparse model parameters to reduce computation and communication overhead. To improve the aggregation phase with sparse models, the works in [17; 20; 21] perform averaging only between overlapping parameters to avoid information dilution by excluding zero value parameters. The authors in [18] obtained a sparse model by selecting a particular client to prune an initial dense model and then performed training in a similar way to FedAvg. In [4; 24], the authors presented binary masks adjustment strategy to improve the performance of sparse models and communication efficiency. The work in [25] progressively pruned a dense model for sparsification and analyzed its convergence. In [19; 22], the clients optimized personalized sparse models by exchanging lottery tickets [26] at every communication round. The work in [5] obtained personalized sparse models by $l_{1}$ norms constraints and the correlation between local and global models. In [8], the authors proposed dual pruning scheme for both local and global models to reduce the communication costs. The FL framework of [23] allows clients to train personalized sparse models in a decentralized setting without a central server. Although these works [17; 18; 4; 24; 25; 19; 22; 5; 23] adopted sparse models during training, they used unstructured pruning, which is difficult to improve the computation efficiency in practice. Meanwhile, with structured sparsity, the authors [7] proposed a training scheme that allows clients to train smaller submodels of a global model. In [9], clients train set of submodels with fixed channel sparsity patterns depending on their computing capabilities. The work in [6] studied structured sparsity by adjusting clients’ channel activation probabilities. However, the works in [7; 9] fixed sparsity patterns and did not optimize sparse model structures. Although [6] optimized channel activation probabilities, the communication cost of downlink still remains high as a server broadcasts whole parameters. Similar to our work, in [27; 28], only binary masks are communicated and optimized by training auxiliary variables to learn sparse model structures. However, the work in [27] approximated binarization step using a sigmoid function during forward propagation. In [28], the downlink communication costs remained the same as that of FedAvg. In [10; 11; 29], clients perform neural-architecture-search by training multiple models to find optimized and sparse models to improve computational and memory efficiency at inference phase. However, in practice, clients often have limited resources to support the computationally intensive architecture search process [30]. Therefore, most prior works either adopted unstructured pruning or they still required extensive computing and communication costs for finding optimal sparse models. In contrast to the prior art, in the proposed SpaFL framework, we find sparse model structures with structured sparsity by optimizing and communicating trainable thresholds for filter/neurons.
@@ 5.2 Baselines
We compare SpaFL with multiple state of the art baselines that studied sparse model structures in FL. In FedAvg [1], every client trains a global dense model and communicates whole model parameters. FedPM [28] trains and communicates a binary mask while freezing model parameters. In HeteroFL [7], each client trains and communicates $p$-reduced models, which remove the last $1-p$ output channels in each layer. In Fjord [9], each client randomly samples a model from a set of $p$-reduced models, which drops out $p\%$ of filter/neurons in each layer. FedP3 [41] communicates a subset of sparse layers that are pruned by the server for downlink and personalize the remaining layers. Clients only upload the updated remaining layers to the server. FedSpa [4] trains personalized sparse models for clients while maintaining fixed model density during training. Local only performs local training with the introduced pruning method without any communications. For the sparse FL baselines, the average target sparsity is set to 0.5 following the configurations in [28; 7; 9; 41; 4].
@@ 5.3 Main Results
In Table LABEL:tab:1 and Fig. 2, we present the averaged accuracy, communication costs, number of FLOPs during training, and convergence rate for each algorithm. We consider all uplink and downlink communications to calculate the communication cost of each algorithm. We also provide the details of the FLOPs measure in the Supplementary document. We average the model densities of SpaFL when a model achieved the best accuracy during training. From these results, we observe that SpaFL outperforms all baselines while using the least amount of communication costs and number of FLOPs. The achieved model densities are 35.36%, 30.57%, and 35.38%, for FMNIST, CIFAR-10, and CIFAR-100, respectively. We also observe that SpaFL uses less resources and performs better than FedP3, HetroFL and Fjord, which deployed structured sparse models across clients. For FedP3, clients only upload subset of layers, but the server still needs to send the remaining layers. Although FedPM reduced uplink communication costs by communicating only binary masks, its downlink cost is the same as FedAvg. In SpaFL, since the clients and the server only exchange thresholds, we can significantly reduce the communication costs compared to baselines that exchange the subset of model parameters such as HeteroFL and Fjord. Moreover, SpaFL significantly achieved better performance than Local, which did not communicate trained thresholds. Local achieved 51.2%, 50.1%, and 53.6% model densities for each dataset, respectively. We can see that communicating trained thresholds can make models sparser and achieve better performance. This also corroborates the analysis of Theorem 1. Hence, SpaFL can efficiently improve model performance with small computation and communication costs. In Fig. 2, we show the convergence rate of each algorithm. We can see that the accuracy of SpaFL decreases and then keeps increasing. The initial accuracy drop is from pruning while global thresholds are not trained enough. As thresholds keep being trained and communicated, clients learn how to prune their model, thereby gradually improving the performance with less active filter/neurons.
@@
Algorithm Accuracy [%] Density [%]
SpaFL 69.78$\pm$2.62 42.2$\pm$4.8
FedAvg 59.20$\pm$0.4 100
Table 4: Performance of SpaFL with the ViT architecture on CIFAR-10
In Tab. 4, we show the performance of SpaFL on a vision transformer using the ViT [42] on CIFAR-10 dataset. We used the same data distribution as done in Tab. LABEL:tab:1. We apply our pruning scheme to multiheads attention layers. Since a multiheads attention layer essentially consists of stacked linear layers, we can simply use (1), thereby making sparse attention. We can see that SpaFL can be applied to transformer architectures by achieving the density of around 42% while outperforming FedAvg.
@@ 6 Conclusion
One limitation of SpaFL is that it cannot explicitly control the sparsity of clients. Since we enforce sparsity through the regularizer term, we need to run multiple experiments to find values for desired sparsity. Another limitation is that our analysis requires a bounded loss function. Meanwhile, in practice, most loss functions may admit bounds that have a large value. For broader impact, SpaFL can reduce not only the computation and communication costs of FL training, but also those of inference phase due to sparsity. Hence, in general, SpaFL can improve the sustainability of FL deployments, and more broadly, of AI.
@@ A.1 Implementation Detail
We run all experiments on NVIDIA A100 GPUs with PyTorch. In Table 5, we provide detailed information of model architectures for each dataset. For the FMNIST dataset, we use the Lenet-5-Caffe model, which is Caffe variant of Lenet-5. The Lenet model has 430500 of model parameters and 580 of trainable thresholds. For the CIFAR-10 dataset, we use a CNN model of seven layers used in Zhou et al. [2019]. It has 807366 of model parameters and 1418 of trainable thresholds. The ResNet-18 model is adopted for the CIFAR-100 dataset with 11159232 of model parameters and 4800 of thresholds. We use a stochastic gradient optimizer with momentum of 0.9. For FMNIST with the Lenet model, we use $\eta(t)=0.001$, $E=5$, a batch size of 64, and $\alpha=0.002$. For CIFAR-10, we use $\eta(t)=0.01$, $E=5$, a batch size of 16, and $\alpha=0.00015$. For CIFAR-100, we use $\eta(t)=0.01$, $E=7$ decayed by $0.993$ at each communication round, a batch size of 64, and $\alpha=0.0007$. All trainable thresholds are initialized to zero. We noticed that too large sparsity coefficient $\alpha$ can dominate the training loss, resulting in masking whole parameters in a certain layer. Following the implementation of Liu et al. [2020], if a certain layer’s density becomes less than 1%, the corresponding trainable thresholds will be reset to zero to avoid masking whole parameters.
For the ViT, we use the patch size of 4, embedding dimension of 128, depth of 6, 8 heads, and set the dimension of linear layers as 256. We use the same setting with the above CIFAR-10 experiments except $\alpha=0.0001$ and $E=1$.
@@ A.1.1 More details about baselines
[…]